Q5.Cheak if ϕ = X2– Y2+ Y represents the velocity potential for 2 dimensional




Irrotational flow. If it does than determine than determine stream function Ψ.

1
Expert's answer
2021-11-01T19:22:45-0400

The condition for irrotational motion


"\dfrac{\partial u}{\partial y}-\dfrac{\partial v}{\partial x}=0"

From the definition of stream function "\u03c8 ," we get "u =\dfrac{\partial \phi}{\partial x}, v=\dfrac{\partial\phi}{\partial y}"

Thus, the velocity components become


"u =\dfrac{\partial \phi}{\partial x}=2x, v=\dfrac{\partial\phi}{\partial y}=-2y+1"

"\dfrac{\partial u}{\partial y}=0, \dfrac{\partial v}{\partial x}=0"


Since

"\dfrac{\partial u}{\partial y}-\dfrac{\partial v}{\partial x}=0-0=0"

then  "\u03d5 = x^2\u2013 y^2+y" represents the velocity potential for 2 dimensional Irrotational flow.



"\dfrac{\partial \psi}{\partial y}=u=2x, -\dfrac{\partial \psi}{\partial x}=v=-2y+1"

Then


"\psi=2xy+g(x)"

"\dfrac{\partial \psi}{\partial x}=2y+g'(x)=-v=2y-1"

"g'(x)=-1"

"g(x)=-x+C"

Stream function is


"\psi=2xy-x+C"